Consider the following transition matri for AllianceCo. Exit 50 .30 .35 Manager Shift Leader .05 .55 .00 Store Associate Store Associate45 Shift Leader Manager .00 .00 .15 65 This year there are 8000 Store Associates, 1000 Shift Leaders, and 700 Managers. What is the predicted availability for Shift Leaders for next year? A.550 O B. 950 C. 1000 D.720
Answer #1

Answer

We have to find the predicted availability, which is same as calculating the expected value

Formula for the expected value is given as

Expected value = (x*P(x))

where x is given values and P(x) is probability given in the above table

setting the given values, we get

Expected value = (8000*0.05) + (1000*0.55) + (700*0.00) = 400 + 550 + 0 = 950

So, required predicted value or expected value is 950